OpenStudy (anonymous):

Is it possible to move an object on the other side of glass with your reflection you see in the glass? since your reflection is made up light and light has energy through the equation E=hv and energy has mass through the equation E=mc^2 and all a force needs is mass and acceleration, is it possible to move an object with a reflection?

OpenStudy (anonymous):

Can you explain the setup of this a little better? I'm not sure how the glass fits into this. This would not be possible on earth because your reflection would not contain enough energy to overcome the other forces acting on the object.

OpenStudy (anonymous):

Yea thats what I was thinking, the glass is maybe a window or something a few feet from you

OpenStudy (anonymous):

I would still say that this is not possible on Earth but in a purely hypothetical situation this would be possible.

OpenStudy (anonymous):

E=mc^2 does not necessarily state that energy has mass, but that mass is a form of energy. This means that if you could take the energy in the light and turn it into a mass, you could possibly use this to have a physical mass to try to hit something in order to move it, but that would mean that the light is no longer light and therefore not reflecting, so you no longer have a reflection there anyway. Moreover, the amount of energy in light is very small compared to the amount of energy in mass (when v=c [speed of light], the energy of a photon would be hc which is on the order of 10^-21J, while the energy in a single gram of matter is .001c^2 which is on the magnitude of 10^14J) so this method would require a lot of light, I believe.

OpenStudy (anonymous):

Your reflection can't move anything on the other side of the mirror, because the light involved never even reaches the other side: What you perceive as your reflection is light that came from whatever light source, hit your body (now it has a color), and was then reflected by the mirror into your eye. At no time is any mass or energy being moved or changed on the other side of the mirror where you perceive the reflection. The photons do have an energy, but they would deposit it onto the mirror itself, not onto anything behind it. Assuming you're wearing a green t-shirt, the light reflected off it would have a wavelength of about 510nm. Its energy would be \[E = h \times v = 6.6 \times 10^{-34} Js \times 510 nm = 3.37 \times 10^{-40} W\] which is an extremely tiny amount. This is also used in the (afaik still fictional) lightsails: A high energy laser pumps light onto a giant reflective surface, the tiny amounts of energy add up and move the mass that is connected to the lightsail forward.

OpenStudy (anonymous):

Well light exerts force if thats what you are asking. If you shine light on something you are essentially pushing it albeit very lightly. This is principal behind solar sails on spaceships. I had a physics teacher who had a tiny wind mill in a vacuum what would spin when it was put under a light. So basically if you are shining light on anything, from a reflection or through glass or anyhow yes, you can move it, but it will be pretty difficult to do just because the force is very small.

OpenStudy (anonymous):

yes